OpenStudy (anonymous):

simplify the radical expression, √20 a. 2√10 b. 5√2 c. 2√5 d. 2

OpenStudy (campbell_st):

\[\sqrt{20} =\sqrt{4 \times 5} = \sqrt{4} \times \sqrt{5 } = 2\sqrt{5}\]